OpenStudy (anonymous):

13. Wind farms are a source of renewable energy found around the world. The power P (in kilowatts) generated by a wind turbine varies directly as the cube of the wind speed v (in meters per second). If a turbine generates 500kW in a 10m/s wind, how much power does it generate in a 12m/s wind?

OpenStudy (anonymous):

power P (in kilowatts) generated by a wind turbine varies directly as the cube of the wind speed v (in meters per second). P=C*v^3, where C = constant, lets find the C first on the given given: P = C*v^3, 500= C(10)^3 what is C ? C=500/(10)^3 ? once you know C plug them to the formula P=C*v^3

OpenStudy (anonymous):

how much power does it generate in a 12m/s wind? since you know C=500/10^3=1/2=0.5 yes or no? plug them to the formula P=C*v^3 P=(0.5)*(12)^3 P= ? ...kilo Watts

OpenStudy (anonymous):

864KW